# 3EED: Ground Everything Everywhere in 3D

### v1

@@ Abstract
Visual grounding in 3D is the key for embodied agents to localize language-referred objects in open-world environments. However, existing benchmarks are limited to indoor focus, single-platform constraints, and small scale. We introduce 3EED, a multi-platform, multi-modal 3D grounding benchmark featuring RGB and LiDAR data from vehicle, drone, and quadruped platforms. We provide over 128,000128{,}000 objects and 22,00022{,}000 validated referring expressions across diverse outdoor scenes – $\mathbf{10\times}$ larger than existing datasets. We develop a scalable annotation pipeline combining vision-language model prompting with human verification to ensure high-quality spatial grounding. To support cross-platform learning, we propose platform-aware normalization and cross-modal alignment techniques, and establish benchmark protocols for in-domain and cross-platform evaluations. Our findings reveal significant performance gaps, highlighting the challenges and opportunities of generalizable 3D grounding. The 3EED dataset and benchmark toolkit are released to advance future research in language-driven 3D embodied perception.
Figure 1: Multi-modal, multi-platform 3D grounding from 3EED. Given a scene and a structured natural language expression, the task is to localize the referred object in 3D space. Our dataset captures diverse embodied viewpoints from Refer to caption Vehicle, Refer to caption Drone, Refer to caption Quadruped platforms, presenting unique challenges in spatial reasoning, scene analysis, and cross-platform 3D generalization.
@@ 1 Introduction
To address these gaps, we introduce 3EED, a large-scale, multi-platform, multi-modal benchmark for 3D visual grounding in outdoor environments (see Fig. 1). Our dataset captures synchronized LiDAR and RGB data from three distinct robotic platforms: [Uncaptioned image] Vehicle, [Uncaptioned image] Drone, [Uncaptioned image] Quadruped. It provides over $\mathbf{128{,}000}$ object instances and $\mathbf{22{,}000}$ human-verified referring expressions, making it $\mathbf{10\times}$ larger than existing outdoor grounding benchmarks, as compared in Tab. 1.
Table 1: Summary of outdoor 3D grounding benchmarks. We compare key features from aspects including: 1Platform ([Uncaptioned image] Vehicle, [Uncaptioned image] Drone, [Uncaptioned image] Quadruped), 2Area Coverage, and 3Statistics. Our dataset exhibits advantages on platform diversity, large collections of LiDAR (L) and camera (C) scenes (Sce.), 3D objects (Obj.), referring expressions (Expr.), and rich elevation variations (Elev.).
@@ 2 Related Work
Existing 3D grounding datasets mainly target small, sensor‐fixed indoor spaces, leaving outdoor, multi-platform scenarios underexplored. To bridge this gap, we curate 3EED, the first 3D grounding dataset that unifies data from [Uncaptioned image] Vehicle, [Uncaptioned image] Drone, [Uncaptioned image] Quadruped platforms. We formalize the multi-modal, multi-platform 3D grounding task in Sec. 3.1, detail a two-stage annotation pipeline in Sec. 3.2, and present statistics that highlight the scale, diversity, and platform balance in Sec. 3.3.
@@ 3.1 Task Formulation: 3D Grounding in the Wild
Figure 3: Examples of multi-platform 3D grounding from the 3EED dataset. There are clear discrepancies across both sensory data (2D & 3D) and referring expressions from the Refer to caption Vehicle, Refer to caption Drone, and Refer to caption Quadruped platforms. For additional examples, kindly refer to the Appendix.
@@ 3.2 Dataset Curation & Annotations
Multi-Platform 3D Data Annotation. We collect [Uncaptioned image] Vehicle sequences from Waymo [76], and [Uncaptioned image] Drone and [Uncaptioned image] Quadruped sequences from M3ED [11]. We adopt a uniform three-stage pipeline for the Drone/Quadruped LiDAR–RGB (see Fig. 2, left). 1) Pseudo-label seeding: State-of-the-art detectors [73, 74, 16, 104, 100, 95] trained on Waymo [76], nuScenes [8], and Lyft [27] produce platform-agnostic 3D boxes for every frame. 2) Automatic consolidation: Kernel-density estimation (KDE) merges detector votes, a 3D multi-object tracker [19] enforces temporal coherence and fills missed detections, and the Tokenize-Anything [63] model is used to project each box onto the RGB view to confirm its class; category conflicts are auto-flagged. 3) Human refinement: Annotators polish the flagged boxes in the user interface, cross-validating to equalize accuracy across platforms. This hybrid scheme yields consistent annotations while limiting manual effort to roughly $100$s per frame.
@@
Referring Expression Data Annotation. After collecting the 3D boxes, we attach platform-invariant language supervision through a parallel procedure (see Fig. 2, middle). 1) Structured prompting: Each 3D box is projected onto its RGB view, together with a knowledge base with five template slots category, status, absolute location, egocentric position, relation, to a vision language model [80]. Few-shot expression examples in the prompt are used to guide the model to output a single, well-formed referring sentence. Platform-specific terms are normalized by platform-invariant rewriting rules to ensure consistent wording across vehicle, drone, and quadruped views. 2) Human verification: Annotators inspect the image, projected box, and caption in an interactive UI, checking semantic correctness, spatial fidelity, absence of ambiguity, and platform-consistency. Cases that are unsatisfactory will be discarded. This staged pipeline delivers concise, unambiguous expressions across vehicle, drone, and quadruped views, providing high-quality language targets for 3D visual grounding.
@@
3.3 Dataset Statistics & Analysis
Benchmark Comparisons. 3EED is, to our knowledge, the first outdoor 3D visual grounding benchmark that standardizes sensing across three embodied platforms [Uncaptioned image]Vehicle, [Uncaptioned image]Drone, and [Uncaptioned image]Quadruped by using synchronized LiDAR–RGB acquisition. As summarized in Tab. 1, our dataset provides $\mathbf{128{,}735}$ object bounding boxes and $\mathbf{22{,}439}$ human-verified referring expressions over $\mathbf{20{,}367}$ tightly time-aligned frames, focusing on the two safety-critical classes Vehicle and Pedestrian. Spatially, our scenes span up to $\mathbf{280\,m\,\times\,240\,m}$ horizontally and exceed $\mathbf{80\,m}$ in elevation, with an order of magnitude larger than any previous outdoor corpus, making it uniquely suited for studying long-range, cross-platform grounding. The train/val split is carefully balanced. As shown in Fig. 4 (middle), containing $2.7\text{k}/2.7\text{k}$ vehicle, $4.1\text{k}/2.9\text{k}$ drone, and $4.9\text{k}/2.9\text{k}$ quadruped scenes, enabling rigorous analysis of both platform-specific challenges and cross-platform generalization.
1) Viewpoint geometry of targets: Fig.4 (left) shows the distribution of pitch angle $\theta^{r}$ and BEV range $\rho$ for each 3D box. [Uncaptioned image] Vehicle data clusters at mid-range with near-zero pitch, typical of level driving. [Uncaptioned image] Drone covers larger $\rho$ with steep negative $\theta^{r}$ from top-down views. [Uncaptioned image] Quadruped stays close in $\rho$ but varies widely in pitch due to ground-level perspective. These patterns expose models to varied spatial cues like “behind” and “under”, improving generalization to novel viewpoints.
2) Per-platform object density: Fig. 4 (middle) shows object density per platform. [Uncaptioned image] Drone captures the busiest scenes due to its wide view, [Uncaptioned image] Vehicle records moderate density, and [Uncaptioned image] Quadruped sees fewer but closer objects. This range enables 3EED to test the ability to disambiguate crowded scenes, maintain situational awareness, and localize small, nearby targets – offering a challenging testbed for robust 3D grounding.
3) Input point-cloud geometry: Fig.4 (right) shows the vertical distribution of LiDAR points $p^{z}$ per platform. [Uncaptioned image] Vehicle scans center around the sensor height, [Uncaptioned image] Drone captures top-down views, and [Uncaptioned image] Quadruped looks upward toward obstacles. These elevation biases affect how spatial terms like “above” or “below” are grounded, offering rich vertical language diversity across viewpoints.
@@ 4.1 Task Suite & Evaluation Strategy
3) Multi-object grounding: A single query may describe multiple referents; the output must localize all targets in the scene. We keep the same IoU thresholds as above but use joint correctness: a query counts as correct iff every referred object is localized correctly. Fig.  5 illustrates several such scenes.
4) Multi-platform grounding: Train on the union of all platforms and evaluate per-platform. We employ balanced design across platforms while keeping the total training budget fixed. This tests whether pooled supervision can close transfer gaps without overfitting to platform-specific statistics.
@@ 5.1 Experimental Setups
Table 4: Ablation study on components. Multi-platform results (Acc@25/50, %) comparing Full vs.removing one module (–CPA, –MSS, –SAF). Dropping any module degrades performance relative to Full, showing their complementarity.
@@ 5.2 Comparative Study
Cross-Platform Generalization. Tab. 2 compares existing 3D grounding backbones under in-distribution (single‐platform) and out‐of‐distribution (cross‐platform) settings.
@@
Coherent Object Co-grounding. Tab. 3 presents the evaluation results on our dataset for the multi-object grounding task. Notably, in this setting, Acc@25 is a strict metric that requires all objects mentioned in the description to be correctly grounded, while mIoU captures the average IoU across individual predicted-ground truth pairs. Existing methods such as BUTD-DETR achieve moderate mIoU ($47.88$) but low joint grounding ($\texttt{Acc@25}=25.40$), revealing their tendency to localize objects in isolation rather than reason about them collectively. In contrast, our baseline leverages multi-scale sampling and dynamic feature fusion to build discriminative representations that capture both fine details and broad context, essential for disambiguating multiple objects of varying size and distance. These design choices deliver substantial improvements in both metrics, demonstrating markedly stronger multi-object reasoning and tighter language-to-3D alignment in complex outdoor scenes.
Qualitative Assessments. Fig. 6 showcases representative multi-platform grounding results on vehicle, drone, and quadruped data. Our unified model consistently outputs precise, tightly aligned 3D boxes despite drastic shifts in viewpoint, object scale, and point-cloud density. In contrast, baseline methods like BUTD-DETR [32] and EDA [87] often yield misaligned or fragmented predictions, especially under challenging aerial and low-angle quadruped perspectives. These comparisons underscore our ability to learn genuine cross-platform invariance and deliver reliable grounding across diverse embodied sensing scenarios.
@@ 5.3 Ablation Study
Component Analysis. Tab. 5 shows that our modules target different sources of error and, together, improve both in-domain accuracy and cross-platform transfer.
@@
Object Density Impact. We analyze how referential grounding performance varies with the object density per scene. We divide test samples into bins based on the number of annotated 3D bounding boxes (11-33, 4466, 7799, 1010+), and compute the average Acc@25 for each bin. As shown in Tab. 5, accuracy consistently drops as object count increases. On the [Uncaptioned image] Quadruped platform, Acc@25 drops from $71.23$ in scenes with 1133 objects to $30.75$ in scenes with 7799 objects. This reflects the increased difficulty of resolving referential ambiguity in cluttered environments.
Platform Complexity Impact. Tab. 6 breaks down grounding performance by platform alongside two key scene statistics: mean LiDAR points per object and mean object count per scene. [Uncaptioned image] Drone scenes suffer the lowest Acc@50, driven by extreme sparsity (just $102$ points/object vs.462462 for [Uncaptioned image] Vehicle and $112$ for [Uncaptioned image] Quadruped) and the highest object density ($8.05$ objects/scene), which together amplify distractors and hinder precise localization. Quadruped data, with moderate density ($112$ points/object) but fewer objects, sits between drone and vehicle performance. These disparities, including ultra-sparse returns and elevated clutter, explain the pronounced aerial performance gap.
@@ 6 Conclusion
Acknowledgments
This study is supported by the National Natural Science Foundation of China (No. 62306257), the Guangzhou Municipal Science and Technology Project (No. 2024A03J0619 and No. 2024A04J4390), and the Guangzhou-HKUST(GZ) Joint Funding Program (Grant No.2023A03J0008), Education Bureau of Guangzhou Municipality.
This study is supported by the Ministry of Education, Singapore, under its MOE AcRF Tier 2 (MOET2EP20221-0012, MOE-T2EP20223-0002). This research is also supported by cash and in-kind funding from NTU S-Lab and industry partner(s).
Lingdong Kong is supported by the Apple Scholars in AI/ML Ph.D. Fellowship program.
The authors would like to sincerely thank the Program Chairs, Area Chairs, and Reviewers for the time and effort devoted during the review process.
Appendix
@@ B.4 Evaluation Metrics
$\text{mIoU}=\frac{1}{N}\sum_{i=1}^{N}\text{IoU}(\hat{b}_{i},b_{i}^{\text{gt}})~.$

### v2

# ] [Uncaptioned image] WorldBench Team

@@ Abstract
Visual grounding in 3D is the key for embodied agents to localize language-referred objects in open-world environments. However, existing benchmarks are limited to indoor focus, single-platform constraints, and small scale. We introduce 3EED, a multi-platform, multi-modal 3D grounding benchmark featuring RGB and LiDAR data from vehicle, drone, and quadruped platforms. We provide over 128,000128{,}000 objects and 22,00022{,}000 validated referring expressions across diverse outdoor scenes – $\mathbf{10\times}$ larger than existing datasets. We develop a scalable annotation pipeline combining vision-language model prompting with human verification to ensure high-quality spatial grounding. To support cross-platform learning, we propose platform-aware normalization and cross-modal alignment techniques, and establish benchmark protocols for in-domain and cross-platform 3D grounding evaluations. Our findings reveal significant performance gaps, highlighting the challenges and opportunities of generalizable 3D grounding. The 3EED dataset and benchmark toolkit have been released to advance future research in language-driven 3D embodied perception.
[…]
[ [Uncaptioned image] Project Page]https://project-3eed.github.io
[…]
Figure 1: Multi-modal, multi-platform 3D grounding from 3EED. Given a scene and a structured natural language expression, the task is to localize the referred object in 3D space. Our dataset captures diverse embodied robot sensing viewpoints from Refer to caption Vehicle, Refer to caption Drone, and Refer to caption Quadruped platforms, presenting unique challenges in spatial reasoning, scene analysis, and cross-platform 3D generalization.
@@ 1 Introduction
To address these gaps, we introduce 3EED, a large-scale, multi-platform, multi-modal benchmark for 3D visual grounding in outdoor environments (see Figure 1). Our dataset captures synchronized LiDAR and RGB data from three distinct robotic platforms: [Uncaptioned image] Vehicle, [Uncaptioned image] Drone, [Uncaptioned image] Quadruped. It provides over $\mathbf{128{,}000}$ object instances and $\mathbf{22{,}000}$ human-verified referring expressions, making it $\mathbf{10\times}$ larger than existing outdoor grounding benchmarks, as compared in Table 1.
Table 1: Summary of existing outdoor 3D visual grounding benchmarks. We compare key features from aspects including: 1Platform ([Uncaptioned image] Vehicle, [Uncaptioned image] Drone, and [Uncaptioned image] Quadruped), 2Area Coverage, and 3Statistics. Our dataset exhibits advantages in platform diversity, large collections of LiDAR (L) and camera (C) scenes (Sce.), 3D objects (Obj.), referring expressions (Expr.), and rich elevation variations (Elev.).
@@ 2 Related Work
Existing 3D grounding datasets mainly target small, sensor‐fixed indoor spaces, leaving outdoor, multi-platform scenarios underexplored. To bridge this gap, we curate 3EED, the first 3D grounding dataset that unifies data from [Uncaptioned image] Vehicle, [Uncaptioned image] Drone, and [Uncaptioned image] Quadruped platforms. We formalize the multi-modal, multi-platform 3D grounding task in Section 3.1, detail a two-stage annotation pipeline in Section 3.2, and present statistics that highlight the scale, diversity, and platform balance in Section 3.3.
@@ 3.2 Dataset Curation & Annotations
Multi-Platform 3D Data Annotation. We collect the [Uncaptioned image] Vehicle sequences from Waymo [76], and [Uncaptioned image] Drone and [Uncaptioned image] Quadruped sequences from M3ED [11]. We adopt a uniform three-stage pipeline for the Drone/Quadruped LiDAR–RGB (see Figure 2, left). 1) Pseudo-label seeding: State-of-the-art detectors [73, 74, 16, 104, 100, 95] trained on Waymo [76], nuScenes [8], and Lyft [27] produce platform-agnostic 3D boxes for every frame. 2) Automatic consolidation: Kernel-density estimation (KDE) merges detector votes, a 3D multi-object tracker [19] enforces temporal coherence and fills missed detections, and the Tokenize-Anything [63] model is used to project each box onto the RGB view to confirm its class; category conflicts are auto-flagged. 3) Human refinement: Annotators polish the flagged boxes in the user interface, cross-validating to equalize accuracy across platforms. This hybrid scheme yields consistent annotations while limiting manual effort to roughly $100$s per frame.
@@
Referring Expression Data Annotation. After collecting the 3D boxes, we attach platform-invariant language supervision through a parallel procedure (see Figure 2, middle). 1) Structured prompting: Each 3D box is projected onto its RGB view, together with a knowledge base with five template slots category, status, absolute location, egocentric position, relation, to a vision language model [80]. Few-shot expression examples in the prompt are used to guide the model to output a single, well-formed referring sentence. Platform-specific terms are normalized by platform-invariant rewriting rules to ensure consistent wording across vehicle, drone, and quadruped views. 2) Human verification: Annotators inspect the image, projected box, and caption in an interactive UI, checking semantic correctness, spatial fidelity, absence of ambiguity, and platform-consistency. Cases that are unsatisfactory will be discarded. This staged pipeline delivers concise, unambiguous expressions across vehicle, drone, and quadruped views, providing high-quality language targets for 3D visual grounding.
Figure 3: Examples of multi-platform 3D grounding from the 3EED dataset. There are clear discrepancies across both sensory data (2D & 3D) and referring expressions from the Refer to caption Vehicle, Refer to caption Drone, and Refer to caption Quadruped platforms.
3.3 Dataset Statistics & Analysis
Benchmark Comparisons. 3EED is, to our knowledge, the first outdoor 3D visual grounding benchmark that standardizes sensing across three embodied platforms [Uncaptioned image]Vehicle, [Uncaptioned image]Drone, and [Uncaptioned image]Quadruped by using synchronized LiDAR–RGB acquisition. As summarized in Table 1, our dataset provides $\mathbf{128{,}735}$ object bounding boxes and $\mathbf{22{,}439}$ human-verified referring expressions over $\mathbf{20{,}367}$ tightly time-aligned frames, focusing on the two safety-critical classes Vehicle and Pedestrian. Spatially, our scenes span up to $\mathbf{280\,m\,\times\,240\,m}$ horizontally and exceed $\mathbf{80\,m}$ in elevation, with an order of magnitude larger than any previous outdoor corpus, making it uniquely suited for studying long-range, cross-platform grounding. The train/val split is carefully balanced. As shown in Figure 4 (middle), containing $2.7\text{k}/2.7\text{k}$ vehicle, $4.1\text{k}/2.9\text{k}$ drone, and $4.9\text{k}/2.9\text{k}$ quadruped scenes, enabling rigorous analysis of both platform-specific challenges and cross-platform generalization.
1) Viewpoint geometry of targets: Figure 4 (left) shows the distribution of pitch angle $\theta^{r}$ and BEV range $\rho$ for each 3D box. [Uncaptioned image] Vehicle data clusters at mid-range with near-zero pitch, typical of level driving. [Uncaptioned image] Drone covers larger $\rho$ with steep negative $\theta^{r}$ from top-down views. [Uncaptioned image] Quadruped stays close in $\rho$ but varies widely in pitch due to ground-level perspective. These patterns expose models to varied spatial cues like “behind” and “under”, improving generalization to novel viewpoints.
2) Per-platform object density: Figure 4 (middle) shows object density per platform. [Uncaptioned image] Drone captures the busiest scenes due to its wide view, [Uncaptioned image] Vehicle records moderate density, and [Uncaptioned image] Quadruped sees fewer but closer objects. This range enables 3EED to test the ability to disambiguate crowded scenes, maintain situational awareness, and localize small, nearby targets – offering a challenging testbed for robust 3D grounding.
3) Input point-cloud geometry: Figure 4 (right) shows the vertical distribution of LiDAR points $p^{z}$ per platform. [Uncaptioned image] Vehicle scans center around the sensor height, [Uncaptioned image] Drone captures top-down views, and [Uncaptioned image] Quadruped looks upward toward obstacles. These elevation biases affect how spatial terms like “above” or “below” are grounded, offering rich vertical language diversity across viewpoints.
@@ 4.1 Task Suite & Evaluation Strategy
3) Multi-object grounding: A single query may describe multiple referents; the output must localize all targets in the scene. We keep the same IoU thresholds as above but use joint correctness: a query counts as correct iff every referred object is localized correctly. Figure 5 illustrates several such scenes.
4) Multi-platform grounding: Train on the union of all platforms and evaluate per-platform. We employ a balanced design across platforms while keeping the total training budget fixed. This tests whether pooled supervision can close transfer gaps without overfitting to platform-specific statistics.
@@ 5.1 Experimental Setups
Table 4: Ablation study on components. Multi-platform results (Acc@25/50, %) comparing Full vs. removing one module (–CPA, –MSS, –SAF). Dropping any module degrades performance, showing their complementarity.
@@ 5.2 Comparative Study
Cross-Platform Generalization. Table 2 compares existing 3D grounding backbones under in-distribution (single‐platform) and out‐of‐distribution (cross‐platform) settings.
@@
Coherent Object Co-grounding. Table 3 presents the evaluation results on our dataset for the multi-object grounding task. Notably, in this setting, Acc@25 is a strict metric that requires all objects mentioned in the description to be correctly grounded, while mIoU captures the average IoU across individual predicted-ground truth pairs. Existing methods such as BUTD-DETR achieve moderate mIoU ($47.88$) but low joint grounding ($\texttt{Acc@25}=25.40$), revealing their tendency to localize objects in isolation rather than reason about them collectively. In contrast, our baseline leverages multi-scale sampling and dynamic feature fusion to build discriminative representations that capture both fine details and broad context, essential for disambiguating multiple objects of varying size and distance. These design choices deliver substantial improvements in both metrics, demonstrating markedly stronger multi-object reasoning and tighter language-to-3D alignment in complex outdoor scenes.
Qualitative Assessments. Figure 6 showcases representative multi-platform grounding results on vehicle, drone, and quadruped data. Our unified model consistently outputs precise, tightly aligned 3D boxes despite drastic shifts in viewpoint, object scale, and point-cloud density. In contrast, baseline methods like BUTD-DETR [32] and EDA [87] often yield misaligned or fragmented predictions, especially under challenging aerial and low-angle quadruped perspectives. These comparisons underscore our ability to learn genuine cross-platform invariance and deliver reliable grounding across diverse embodied sensing scenarios.
@@ 5.3 Ablation Study
Component Analysis. Table 5 shows that our modules target different sources of error and, together, improve both in-domain accuracy and cross-platform transfer.
@@
Object Density Impact. We analyze how referential grounding performance varies with the object density per scene. We divide test samples into bins based on the number of annotated 3D bounding boxes (11-33, 4466, 7799, 1010+), and compute the average Acc@25 for each bin. As shown in Table 5, accuracy consistently drops as object count increases. On the [Uncaptioned image] Quadruped platform, Acc@25 drops from $71.23$ in scenes with 1133 objects to $30.75$ in scenes with 7799 objects. This reflects the increased difficulty of resolving referential ambiguity in cluttered environments.
Platform Complexity Impact. Table 6 breaks down grounding performance by platform alongside two key scene statistics: mean LiDAR points per object and mean object count per scene. [Uncaptioned image] Drone scenes suffer the lowest Acc@50, driven by extreme sparsity (just $102$ points/object vs.462462 for [Uncaptioned image] Vehicle and $112$ for [Uncaptioned image] Quadruped) and the highest object density ($8.05$ objects/scene), which together amplify distractors and hinder precise localization. Quadruped data, with moderate density ($112$ points/object) but fewer objects, sits between drone and vehicle performance. These disparities, including ultra-sparse returns and elevated clutter, explain the pronounced aerial performance gap.
@@ 6 Conclusion
[…]
@@ B.4 Evaluation Metrics
$\text{mIoU}=\frac{1}{N}\sum_{i=1}^{N}\text{IoU}(\hat{b}_{i},b_{i}^{\text{gt}})\penalty 10000\ .$